Make a study about how many sheets of paper you consumed weekly in answering




your Self Learning Modules. Record the quantity (total number of sheets) per subject, then




construct a probability distribution. Compute the mean, variance, and the standard deviation




of the probability distribution you made. Interpret the result, then find out how many weeks




you will consume 50 sheets of pad paper.





1
Expert's answer
2022-02-22T23:02:46-0500

we define the following as my observations

number of subjects(x) p(x)

1 0.15

2 0.28

3 0.14

4 0.26

5 0.17



Mean number of sheets "\mu = \sum X. P(X)=0.15+0.56+0.42+1.04+0.85=3.02"


Variance "\sigma^2=" "\sum[X^2.P(X)]-\mu^2= [0.15+1.12+1.26+4.16+4.25]-3.02"

"=10.94-3.02=7.92"


Standard Deviation "\sigma = \sqrt{7.92}=2.81"


So, no. of weeks required to use 50 sheets of paper = "\dfrac{50}{3.02}=16.56 \approx 17 \ weeks"